OpenStudy (anonymous):

factor -x^2+25x-625=0

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

then what can it be solved for?

OpenStudy (a_clan):

\[x ^{2} - Sx +P=0\] where S= a + b and P = a.b a and b are the two roots of quadratic equation.

OpenStudy (anonymous):

factor x/25 = 25/25x